OpenStudy (anonymous):

If P(A or B) = 0.8, P(A) = 0.1 and P(B) = 0.9, determine P(A and B). 0.90 0.20 0.80 0.08

OpenStudy (jim766):

P(A or B)= P(A) + P(B) - P(A and B) 0.8=0.1+0.9-x solve for x

OpenStudy (anonymous):

would x be .80

OpenStudy (jim766):

try that again..... .8 = .1 + .9 - x .8 = .10 - x

OpenStudy (jim766):

how do you get the x alone?

OpenStudy (jim766):

what is keeping the x from being by itself?

OpenStudy (anonymous):

i really dont undertand this math

OpenStudy (jim766):

so forget about the math....in the equation .8 = .10 - x what is keepng the x from being by itself?

OpenStudy (anonymous):

what is x

OpenStudy (jim766):

.8 = .10 - x -.10 -.10 -.2 = -x .2 = x